OpenStudy (crashonce):

A committe of five is to b chosen from six men and eight women. Find how many committees are possible if a particular man must be included? The answer is715 but i don't understand. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

"A committe of five ... a particular man must be included" so you really have 4 slots left because this particular man must be locked in.

jimthompson5910 (jim_thompson5910):

6 men total 6-1 = 5 men left over after that person is locked in 8 women total so there are 8+5 = 13 people who aren't locked in there are 4 slots, so you need to compute C(13,4)

jimthompson5910 (jim_thompson5910):

Formula: \[\Large C(n,r) = \frac{n!}{r!(n-r)!}\]

OpenStudy (crashonce):

thanks so much

jimthompson5910 (jim_thompson5910):

np